Arif took a loan of ₹ 80,000 from a bank. If the rate of interest is 10% per annum, find the difference in amounts he would be paying after 1whole 𝟏/𝟐 years if the interest is 
Compounded annually 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

P =₹80,000
R = 10% per annum
n = 1 whole 1/2years.
The amount for 1 year and 6 months can be calculated by first calculating the amount for
1 year using the compound interest formula, and then calculating the simple interest for 6 months on the amount obtained at the end of 1 year.
First, the amount for 1 year has to be calculated.
Amount, A =P (1 +R/100 )^n
=₹ [8000(1+ 10/100 )^1]
=₹ [8000 × (11/100)^3]
=₹88000
By taking₹88,000 as principal, the SI for the next1/2year will be calculated.
S.I. = P x R x T
100
= 88000 x 10 x 1/2
100
=₹4400
Interest for the first year =₹ 88000 − ₹80000 =₹8,000
And interest for the next1/2year =₹4,400
Total C.I. =₹8000 +₹4,400 =₹1,2400
A = P + C.I. =₹(80000 + 12400)
=₹92,400

​Thanks